Find the area under the standard normal distribution curve.



1. Between z = 0 and z = 1.66



2. Between z = 0 and z = -0.35



3. To the right of z = 1.10



4. To the left of z = -0.48



5. Between z = 1.23 and z = 1.90



6. Between z = - 1.56 and z = -1.83



7. To the left of z = - 2.15 and to the right of z = 1.62



8. To the left of z = 2.11



9. To the right of z = -0.15



10. To the right of z = 1.92 and to the left of z = 0.44

1
Expert's answer
2022-01-20T11:38:31-0500

1)P(0<z<1.66)=p(z<1.66)-p(z<0)

=0.9515-0.5=0.4515

2)p(0<z<-0.35)=p(z<-0.35)-p(z<0)

=0.6368-0.5 = 0.1368

3)p(z>1.10)=1-p(z<1.10)

=1-0.8643=0.1357

4)p(z<-0.48)=1-p(z>-0.48)

=1-0.3156 = 0.6844

5)p(1.23<z<1.90)=p(z<1.90)-p(z<1.23)

=0.9713-0.1093=0.862

6)p(-1.83<z<-1.56)=p(z<-1.56)-p(z<-1.83)

0.0594-0.0336=0.0258

7)p(z<-2.15)=1-p(z>-2.15)

=1- 0.1578=0.8422.

p(z>1.62)=1-p(z<1.62)

=1- 0.9474= 0.0526

8)p(z<2.11)=1-p(z>2.11)

=1- 0.0175=0.9825

9)p(z>-0.15)=1-p(z<-0.15)

=1- 0.5596= 0.4404

10)p(z>1.92)=1-p(z<1.92)

=1- 0.0274= 0.9726